OpenStudy (anonymous):

Rewrite y = x^2 + 14x + 29 in general form please show your work

OpenStudy (anonymous):

what the monkey is general form? looks good as it is.

OpenStudy (anonymous):

Do you mean like: -x^2 - 14x + y - 29 = 0

OpenStudy (anonymous):

You mean x^2+14x+29-y=0 If yes, you just have to transfer the y and add the - ;)

OpenStudy (anonymous):

\[ y = x^2 + 14x + 29\] This is already in general form.

OpenStudy (anonymous):

\[x^2+14x+29-y=0\] Here it is.